OpenStudy (anonymous):

series (n!)^n / n^4n converges or diverges?

OpenStudy (anonymous):

try nth root test. i believe it diverges.

OpenStudy (anonymous):

If you just neet to check for convergence or divergence, why not just put the values and check, \[1 \rightarrow 1!^1/1^4 = 1\] \[2 \rightarrow 2!^2/2^8 = 4/256 = 1/64\]

OpenStudy (anonymous):

matter of fact these terms don't even go to zero, so it clearly diverges. as an example \[\frac{(10!)^{10}}{10^{40}}\] is huge. the denominator is \[10^{40}\] while the numerator is about \[3.96\times 10^{65}\]

OpenStudy (anonymous):

First tell me, does converging only mean that the terms go on becoming small and small?

OpenStudy (anonymous):

If so, (n+1)!^(n+1)/(n+1)^4(n+1) =(n+1)^(n+1)(n!)^(n+1) / (n+1)^4(n+1) \[=(n!)^n (n+1)^{n+1} n! / (n+1)^{(n+1)^{4}}\]

OpenStudy (anonymous):

So it has to converge!